OpenStudy (anonymous):

The perimeter of a rectangle is 10 cm, and its dimensions are tripled. What is the perimeter of the larger rectangle?

OpenStudy (king):

p=a+b+c new p=3a+3b+3c =3(a+b+c) =3p